# CogDDN: A Cognitive Demand-Driven Navigation with Decision Optimization and Dual-Process Thinking

### v1

@@ 1. Introduction
Mobile robots are increasingly essential for assisting humans in unknown and unstructured environments, such as households, hospitals, and warehouses. To operate effectively, they must be capable of interpreting and responding to implicit human demands, even when target objects are uncertain or not explicitly specified. For instance, a hungry person instinctively seeks food based on internal cues rather than direct instructions. Similarly, demand-driven navigation (DDN) (Wang et al., 2023, 2024c) enables robots to locate objects that fulfill such implicit needs without relying on predefined object lists or locations. However, as shown in Figure 1 (a), conventional data-driven DDN methods (Wang et al., 2023, 2024c) depend heavily on extensive pre-collected data for model training and decision-making, limiting their generalization to novel scenarios and vague instructions.
@@
Recent breakthroughs in Large Language Models (LLMs) and Vision-Language Models (VLMs) have demonstrated strong reasoning capabilities and extensive world knowledge, making them well-suited for navigation tasks. Cognition-driven approaches (Yin et al., 2024; Nie et al., 2025; Zhang et al., 2025; Wang et al., 2024b; Shah et al., 2022) leverage these models to help agents interpret instructions and locate target objects. The DDN tasks combine natural language with visual perception, requiring agents to infer and act upon implicit user intent. As illustrated in Figure 1 (b), our framework integrates LLMs and VLMs for DDN, enabling more sophisticated reasoning over vague or ambiguous instructions. Additionally, VLMs improve navigation by jointly processing visual and linguistic inputs, effectively bridging semantic intent with spatial context. This cognition-driven, multi-agent approach enhances the agent’s spatial and contextual understanding, resulting in more accurate and adaptable navigation. Figure 1 (c) shows that the cognition-driven method generalizes better than traditional data-driven approaches, exhibiting a smaller performance drop between seen and unseen scenarios.
@@
This paper presents CogDDN, a novel VLM-based approach inspired by the dual-process framework for closed-loop decision-making. A key distinction is that it emulates human cognitive and learning mechanisms. Unlike many VLM-based navigation systems (Long et al., 2024; Cai et al., 2024) that rely on additional views, depth, or maps, our system uses only front-facing views, aligning with human search behavior and real-world robotic constraints. At a result, our system performs a more realistic navigation task by mirroring the perceptual constraints and decision-making processes of human object search.
In parallel with human decision-making, CogDDN incorporates a demand-matching module that leverages LLMs to accurately link instructions to relevant objects, allowing the agent to identify and locate key items that satisfy implicit user needs. Furthermore, we introduce a dual-process decision-making system based on dual-process theory (Kahneman, 2011; Evans and Stanovich, 2013; Wason and Evans, 1974), which simulates human cognition through two components: a Heuristic Process (System-I) and an Analytic Process (System-II) (Mei et al., 2024; Ma et al., 2025). The Heuristic Process enables rapid, intuitive decisions based on prior knowledge, while the Analytic Process refines these decisions by analyzing new data and updating the knowledge base. Before navigation, we construct a knowledge base to store high-quality strategies adaptable to diverse scenarios. The Heuristic Process utilizes this accumulated knowledge, refined through supervised fine-tuning (SFT), to guide initial decisions. Meanwhile, the Analytic Process updates the knowledge base when encountering obstacles or challenges, enabling continuous learning and improved decision-making. In contrast to direct end-to-end methods, our system integrates Chain of Thought (CoT) (Wei et al., 2022) reasoning, enhancing the agent’s ability to reason through complex scenarios and make informed decisions, rather than merely outputting final actions.
The main contributions are summarized as follows:
We propose CogDDN, a novel VLM-based framework for DDN tasks that simulates human-like cognitive mechanisms and learning processes.
We introduce a dual-process decision-making module. Without human involvement, our approach allows the empirical Heuristic Process to inherit the capabilities of the rational Analytic Process through a self-supervised mechanism.
CogDDN uses the Analytic Process and a knowledge accumulation mechanism to enrich a transferable knowledge base, enabling continual learning and effective generalization to novel navigation environments.
@@ 2.1. Demand-driven Navigation
Demand-driven Navigation (DDN) (Wang et al., 2023, 2024c) is an innovative task requiring an agent to navigate an environment based on high-level human demands described in natural language. Unlike traditional visual navigation tasks like Visual Object Navigation (VON) (Ramakrishnan et al., 2022; Zhao et al., 2023; Kiran et al., 2022; Chaplot et al., 2020; Gadre et al., 2022; Pal et al., 2021) and Visual Language Navigation (VLN) (Wu et al., 2024; Yue et al., 2024; Zhang et al., 2024b; Zhou et al., 2024a, b; Hong et al., 2025; Goetting et al., 2024; Cai et al., 2024), DDN does not focus on specific object categories or step-by-step instructions but rather on fulfilling user-defined needs. Using abstract, high-level instructions, DDN enables the agent to infer which objects in the environment are relevant to the given demand. This makes DDN more flexible and dynamic than VON or VLN, as it centers on addressing ”human demands” rather than specific objects. The agent must interpret these demands and make inferences from the current scene, providing a more generalized approach to navigation. The main challenge of DDN lies in interpreting and fulfilling human demand within the visual context.
@@ 2.2. Transitioning from Data-Driven to Cognition-Driven Navigation
Early navigation methods were predominantly data-driven, relying on imitation learning (Nguyen et al., 2019; Kamath et al., 2023) and reinforcement learning (Wang et al., 2018; Chen et al., 2022) to train reactive policies on large-scale labeled datasets. While these approaches achieved promising results within their training distributions, they often failed to generalize to out-of-distribution scenarios (Cai et al., 2025; Zhan et al., 2024), mainly due to limited semantic reasoning and contextual awareness, and often incur high computational costs during both training and deployment (Zhang et al., 2024c; Wei et al., 2024; Xiang and Zhang, 2024).
Motivated by the success of pre-trained vision-language models in various multimodal reasoning tasks (Ma et al., 2024; Wang et al., 2024d; Lu et al., 2025; Liu et al., 2025a, b), recent navigation research has begun to adopt these models as cognitive priors to move toward more robust and human-like navigation strategies (Yin et al., 2024; Nie et al., 2025; Zhang et al., 2025). Inspired by human-like reasoning, these systems unify perception and language to support language-driven goal specification (Mirjalili et al., 2024) and semantic understanding (Bangguo Yu and Cao, 2024). Unlike traditional methods that depend on structured instructions or spatial maps, VLM-based approaches allow agents to interpret open-ended natural language commands and navigate in complex, unfamiliar environments (Zhang et al., 2024a; Song et al., 2025). These systems emphasize multimodal grounding, context-aware planning, and continual learning, and they often incorporate reasoning techniques such as Chain-of-Thought to support interpretable and generalizable behavior.
@@ 3.1. Task Definition
CogDDN tasks require agents to locate relevant objects in a given environment based on human demands and perceptual inputs. They can be formalized as a Partially Observable Markov Decision Process (POMDP), defined as
where represents the state space (the agent’s current state), is the observation space (including human demands and images), is the action space (e.g., MoveAhead, RotateLeft, LookDown, etc.), and is the action execution function. For instance, a user may state, ”I am thirsty.” In response, the agent must identify an object capable of quenching thirst based on the current visual observation and then execute the actions to locate the object. After each action, the agent receives updated observations and generates new actions based on the updated information to continue searching for the object. The task is complete once the agent is within 1.5 meters of the target object.
@@ 3.2. Demand Matching
Objects that fulfill the same human demand in demand-driven navigation share key properties. For instance, paintings, houseplants, and statues are suitable candidates for decorating a space, as they enhance visual appeal and complement the intended environment. This relationship between human demands and object properties is grounded in universal knowledge.
While LLMs are proficient in reasoning based on instructions and properties, they encounter difficulties when an exact match for a request is unavailable. In such cases, LLMs may suggest suboptimal objects. For instance, if a user requests, ”I need something to hold my flowers,” a mug might be suggested despite being a less-than-ideal choice. This tendency significantly reduces the model’s accuracy in selecting the most suitable object. To mitigate this issue, we apply supervised fine-tuning (SFT) to train the LLM, allowing it better to align objects with user requirements. This method improves the model’s capacity to identify and recommend the most appropriate objects, ensuring it avoids suggesting less optimal options, even when an exact match is not readily apparent.
Formally, the Demand Matching module implements a target objects generation function , where represents the space of observation space including human demands and images, is the space of demand properties, and is the space of target objects. By leveraging a fine-tuned LLM for demand matching, the system benefits from its ability to reason over complex instructions and object properties. This capability prevents it from forcing matches between human demands and objects when an exact match is unavailable, ensuring more appropriate suggestions even in ambiguous scenarios. Please refer to Appendix A.2 for annotation details.
@@ 3.3. Knowledge Base
Inspired by LeapVAD (Ma et al., 2025), we generate additional data with outputs divided into three components: 1) Scene Description: This component outlines the environment based on the current viewpoint image, highlighting passable areas and task-related objects that may affect mobility. 2) Reasoning: This component leverages the visual-language model’s spatial awareness and commonsense reasoning capabilities based on the target’s location and the scene description to guide the agent’s navigation. 3) Decision: Drawing from the output of the reasoning component, the model generates the optimal decision for the current navigation task, ensuring the selection of the most effective action. This process incorporates knowledge transfer mechanisms to enhance navigation performance.
Building on this foundational approach of leveraging scene descriptions and reasoning for navigation optimization, we construct a knowledge base that integrates environmental data with dynamic decision-making processes. As illustrated in Figure 4, during the execution of the trajectory generated by the A* algorithm (Hart et al., 1968), once the target object is detected, the current viewpoint image , target object , and final action S are input into the VLM to generate scene descriptions D and reasoning R, which are then recorded as an experience in the knowledge base. This accumulated experience can be progressively transferred to the Heuristic Process. As detailed in Section 3.4, we utilize the collected data {} to train the Heuristic Process’s ability to follow instructions and adhere to formatting standards. Please refer to Figure 9 in Appendix A.3 for the data format.
@@ 3.4. Heuristic Process
While the VLM effectively generates reasoning based on decisions from the A* algorithm (Hart et al., 1968), it struggles to directly infer the relationship between reasoning and decisions. Moreover, its relatively slow processing speed often leads to redundancy and repetition, limiting practical use in indoor navigation. Formally, the Heuristic Process module implements an action execution function , where represents the space of target objects, depicts the state space, and is the action space, which is presented in table 1. Inspired by human navigation, where efficiency is gained through repeated practice and environmental familiarity, allowing individuals to navigate with minimal cognitive effort, we introduce the Heuristic Process in CogDDN. As demonstrated in Figure 5, it consists of two modules: Explore and Exploit, both leveraging Chain-of-Thought (CoT) reasoning to adapt to new environments and optimize learned navigation execution. The Heuristic Process is expressed as:
Explore. The Explore module is activated when the Demand Matching module cannot identify a suitable target within the current scene. In such cases, the system searches the environment by generating exploratory actions based on its visual perspective, utilizing the VLM. This aims to uncover previously overlooked objects or paths. As Figure 2 shows, the system first generates a Scene Description D based on the current viewpoint image . Subsequently, it conducts Reasoning R to direct the exploration process. The system’s Decision S consists of a series of actions to explore unvisited regions. The module incorporates information from prior actions and recent rotations to minimize redundant exploration. If obstacles hinder progress, the flag is set to true, prompting the system to dynamically generate new exploratory actions, ensuring continuous progress without unnecessary repetition or resource depletion.
@@
Exploit. Once the Demand Matching module identifies a target object, the system transitions to the Exploit module. Unlike the Explore phase, which focuses on broader navigation and search strategies, the Exploit phase emphasizes precise, deliberate actions to achieve the navigation goal. This phase leverages a fine-tuned version of the VLM, optimized through prior exploration and experiences stored in the knowledge base. Fine-tuning improves the model’s ability to make accurate predictions and decisions based on subtle environmental cues, enabling the system to handle complex tasks with high precision. Guided by CoT, the system first generates a Scene Description D based on the current viewpoint image and the information of target objects , performs Reasoning R to infer the most suitable course of action based on the D and , and produces the most optimal Decision S, a single action focused on reaching the target. Focusing on one action rather than multiple potential actions improves efficiency and computational performance, ensuring rapid and effective goal achievement.

### v2

@@ 1. Introduction
Mobile robots are becoming indispensable in dynamic and unstructured environments such as households, hospitals, and warehouses. To operate effectively and naturally within these human-centric spaces, they must move beyond simple command execution to interpret and respond to implicit human demands. This capability is particularly critical when target objects are uncertain or not explicitly specified. For instance, a hungry person instinctively seeks food based on internal cues, rather than following a direct instruction like ”find the apple on the kitchen counter”. Similarly, demand-driven navigation (DDN) (Wang et al., 2023, 2024c) enables robots to locate objects that fulfill such implicit needs without relying on predefined object lists or known locations. However, as illustrated in Figure 1 (a), conventional data-driven DDN methods (Wang et al., 2023, 2024c) are fundamentally constrained by their heavy reliance on extensive, pre-collected datasets for training and decision-making. This dependence curtails their adaptability and generalization when encountering novel scenarios or vaguely phrased instructions.
@@
Recent breakthroughs in Large Language Models (LLMs) and Vision-Language Models (VLMs) have catalyzed a new era in robotic intelligence, as these models possess remarkable reasoning capabilities and extensive world knowledge that are exceptionally well-suited for complex navigation tasks. Leveraging these advancements, cognition-driven approaches (Yin et al., 2024; Nie et al., 2025; Zhang et al., 2025; Wang et al., 2024b; Shah et al., 2022) employ these models to help agents interpret high-level instructions and strategically locate targets. The DDN task, which inherently fuses natural language with real-time visual perception, demands that an agent infer and act upon subtle user intent. As depicted in Figure 1 (b), our proposed framework synergistically integrates LLMs and VLMs for DDN, enabling sophisticated and nuanced reasoning over ambiguous instructions. The embedded VLMs enhance navigation by jointly processing visual and linguistic inputs, effectively bridging the semantic gap between high-level intent and spatial context. This cognition-driven, multi-agent approach significantly improves the agent’s spatial and contextual understanding, yielding more accurate, robust, and adaptable navigation. Figure 1 (c) shows that the cognition-driven method generalizes better than traditional data-driven approaches, exhibiting a substantially smaller performance drop when transitioning from seen to unseen scenarios.
@@
This paper introduces CogDDN, a novel VLM-based framework for DDN that is inspired by the human dual-process cognitive framework and designed for robust, closed-loop decision-making. A key differentiator of our approach is its emulation of human cognitive and learning mechanisms to solve complex search tasks. Unlike many contemporary VLM-based navigation systems (Long et al., 2024; Cai et al., 2024) that rely on supplementary data such as multiple camera views, depth information, or pre-built maps, our system uses only egocentric, front-facing visual input. This design choice intentionally aligns our agent’s perception with natural human search behavior and adheres to the practical constraints of real-world robotic hardware. Consequently, our system addresses a more realistic and challenging navigation task by mirroring the perceptual limitations and sequential decision-making inherent in human object search.
Analogous to human cognitive processes, CogDDN incorporates a demand-matching module that leverages the semantic reasoning power of LLMs. This module translates ambiguous natural language instructions into a set of relevant, potential target objects, enabling the agent to identify and prioritize items that satisfy the user’s underlying needs. Furthermore, we introduce a core decision-making system founded on dual-process theory (Kahneman, 2011; Evans and Stanovich, 2013; Wason and Evans, 1974), which simulates human cognition via two distinct yet complementary components: a Heuristic Process (System-I) and an Analytic Process (System-II) (Mei et al., 2024; Ma et al., 2025). The Heuristic Process, akin to human intuition, facilitates rapid and efficient decisions based on prior knowledge. In contrast, the Analytic Process employs deliberate, step-by-step reasoning to refine these decisions, particularly in novel or challenging situations, and updates the knowledge base accordingly. Prior to navigation, we construct a knowledge base of high-quality, generalizable strategies. The Heuristic Process, continuously refined through supervised fine-tuning (SFT), draws upon this accumulated knowledge to guide initial actions. Meanwhile, the Analytic Process is invoked to update this knowledge base upon encountering obstacles or failures, creating a virtuous cycle of continuous learning and performance improvement. In contrast to direct end-to-end methods, our system integrates Chain of Thought (CoT) (Wei et al., 2022) reasoning, enhancing the agent’s ability to logically reason through complex scenarios and make transparent, informed decisions, rather than merely outputting final actions.
The main contributions of this paper are summarized as follows:
We propose CogDDN, a novel VLM-based framework for DDN tasks that simulates human-like cognitive mechanisms and learning processes to navigate based on implicit demands.
We introduce a dual-process decision-making module wherein the empirical Heuristic Process learns from the rational Analytic Process via a self-supervised mechanism, eliminating the need for human intervention.
CogDDN uses the Analytic Process and a knowledge accumulation mechanism to progressively enrich a transferable knowledge base, enabling continual learning and effective generalization to novel navigation environments.
@@ 2.1. Demand-driven Navigation
Demand-driven Navigation (DDN) (Wang et al., 2023, 2024c) represents an advanced navigation paradigm where an agent must interpret and act upon high-level human demands expressed in natural language. This task diverges significantly from traditional visual navigation frameworks such as Visual Object Navigation (VON) (Ramakrishnan et al., 2022; Zhao et al., 2023; Kiran et al., 2022; Chaplot et al., 2020; Gadre et al., 2022; Pal et al., 2021), which typically requires finding a specific object category (e.g., ”find a chair”), and Visual Language Navigation (VLN) (Wu et al., 2024; Yue et al., 2024; Zhang et al., 2024b; Zhou et al., 2024a, b; Hong et al., 2025; Goetting et al., 2024; Cai et al., 2024), which relies on explicit, step-by-step instructions (e.g., ”walk past the table and turn left”). In contrast, DDN centers on fulfilling an underlying user need (e.g., ”I’m thirsty”), compelling the agent to infer which objects in the environment are relevant to satisfying the stated demand. This formulation makes DDN a more flexible and dynamic, shifting the focus from locating specific objects to addressing abstract human demands. The agent is thus required to perform a higher level of semantic interpretation and contextual inference based on the current visual scene, fostering a more generalized and human-aligned approach to navigation. Therefore, the primary challenge in DDN lies in this sophisticated process of interpreting abstract human needs and grounding them within a rich, dynamic visual context to guide action.
@@ 2.2. Transitioning from Data-Driven to Cognition-Driven Navigation
Early approaches to robotic navigation were predominantly data-driven, relying heavily on imitation learning (Nguyen et al., 2019; Kamath et al., 2023) and reinforcement learning (Wang et al., 2018; Chen et al., 2022) to train reactive policies. These methods learn to map sensory inputs directly to actions by training on large-scale, often manually labeled, datasets. While such approaches have demonstrated promising results within their training distributions, they frequently exhibit brittleness and fail to generalize to out-of-distribution scenarios (Cai et al., 2025; Zhan et al., 2024). This limitation arises primarily from a lack of deep semantic reasoning and contextual awareness, as the models often learn spurious correlations in the data rather than robust, causal relationships. Furthermore, these data-intensive methods typically incur high computational costs during extensive training and deployment (Zhang et al., 2024c; Wei et al., 2024; Xiang and Zhang, 2024).
Motivated by the success of pre-trained vision-language models in a wide array of multi-modal reasoning tasks (Ma et al., 2024; Wang et al., 2024d; Lu et al., 2025; Liu et al., 2025a, b), recent navigation research has shifted towards cognition-driven paradigms. This new wave of research adopts large-scale models as powerful cognitive priors, aiming to instill more robust and human-like navigation strategies (Yin et al., 2024; Nie et al., 2025; Zhang et al., 2025). Inspired by human reasoning, these systems unify perception, action, and language to support flexible, language-driven goal specification (Mirjalili et al., 2024) and a deeper semantic understanding of the world (Bangguo Yu and Cao, 2024). Unlike traditional methods that depend on structured instructions or explicit spatial maps, VLM-based approaches empower agents to interpret open-ended natural language commands and navigate effectively in complex, unfamiliar environments (Zhang et al., 2024a; Song et al., 2025). These modern systems emphasize multi-modal grounding, context-aware planning, and continual learning. To enhance transparency and robustness, they often incorporate explicit reasoning techniques, such as Chain-of-Thought (CoT) prompting, which encourages the model to generate a sequence of logical steps, thereby supporting more interpretable and generalizable navigation behavior.
@@ 3.1. Task Definition
CogDDN tasks require agents to locate the relevant objects in a given environment based on human demands and perceptual inputs. These tasks can be formally described as a Partially Observable Markov Decision Process (POMDP), defined as
where represents the state space (the agent’s current state), is the observation space (including human demands and images), is the action space (e.g., MoveAhead, RotateLeft, LookDown, etc.), and is the action execution function. For instance, a user may state, ”I am thirsty”. In response, the agent must identify an object capable of quenching thirst based on the current visual observation and then execute the actions to locate the object. After each action, the agent receives updated observations and generates new actions based on the updated information to continue searching for the object. The task is considered complete when the distance between the agent and the target object is reduced to within 1.5 meters.
@@ 3.2. Demand Matching
A foundational principle of demand-driven navigation is that objects capable of fulfilling the same high-level human demand often share a set of key, underlying properties. For instance, paintings, houseplants, and statues are all suitable candidates for decorating a space. Although physically distinct, they all enhance visual appeal and complement the ambient environment. This intrinsic relationship between abstract human demands and object properties is grounded in commonsense or universal knowledge.
While LLMs are proficient at general reasoning based on instructions and properties, they can encounter difficulties when a perfect match for a given request is unavailable in the immediate environment. In such cases, standard LLMs may suggest suboptimal or functionally adjacent objects. For example, if a user requests, ”I need something to hold my flowers”, and no vase is present, the model might suggest a mug. Although a mug can technically hold water and flowers, it is a less-than-ideal choice domestically. This tendency to over-generalize based on basic affordances can significantly reduce the model’s accuracy in selecting the most suitable object that aligns with user expectations. To mitigate this issue, we employ supervised fine-tuning (SFT) to further train the LLM, allowing it to better align object affordances with nuanced user requirements and contextual appropriateness. The fine-tuning enhances the model’s capacity to identify and recommend the most fitting objects, ensuring it avoids suggesting less optimal options, even when an exact match is not readily apparent.
Formally, the Demand Matching module implements a target objects generation function , where represents the space of observation space including the human demands and visual information from images, is the latent space of inferred demand properties, and is the resulting space of candidate target objects. By leveraging a fine-tuned LLM for demand matching, the system benefits from an enhanced ability to reason over complex instructions and subtle object properties. This specialized capability prevents it from forcing matches between human demands and available objects when a direct correspondence is absent, ensuring more appropriate suggestions even in ambiguous scenarios. Please refer to Appendix A.2 for annotation details.
@@ 3.3. Knowledge Base
Inspired by LeapVAD (Ma et al., 2025), we generate additional data with outputs divided into three components: 1) Scene Description: This component outlines the environment based on the current viewpoint image, highlighting passable areas and task-related objects that may affect mobility. 2) Reasoning: This component leverages the visual-language model’s spatial awareness and commonsense reasoning capabilities based on the target’s location and the scene description to guide the agent’s navigation. 3) Decision: Drawing from the output of the reasoning component, the model generates the optimal decision for the current navigation task, ensuring the most effective action. This process incorporates knowledge transfer mechanisms to enhance navigation performance.
Building on this foundational approach of leveraging scene descriptions and reasoning for navigation optimization, we construct a knowledge base that integrates environmental data with dynamic decision-making processes. As illustrated in Figure 4, during the execution of the trajectory generated by the A* algorithm (Hart et al., 1968), once the target object is detected, the current viewpoint image , target object , and final action S are input into the VLMs to generate scene descriptions D and reasoning R, which are then recorded as an experience in the knowledge base. This accumulated experience can be progressively transferred to the Heuristic Process. As detailed in Section 3.4, we utilize the collected data {} to train the Heuristic Process’s ability to follow instructions and adhere to formatting standards. Please refer to Figure 9 in Appendix A.3 for the data format.
@@ 3.4. Heuristic Process
While the VLMs can effectively generate reasoning that aligns with decisions from path planning algorithms like A* algorithm (Hart et al., 1968), they often struggle to directly infer the causal relationship between reasoning and decisions. Moreover, their relatively slow processing speed can lead to redundancy and repetition, limiting their practical application in real-time indoor navigation. To address these challenges, we introduce the Heuristic Process in CogDDN. This process is inspired by human navigation, where efficiency is gained through repeated practice and environmental familiarity, eventually allowing individuals to navigate with minimal cognitive effort. Formally, the Heuristic Process module implements an action execution function , where represents the space of target objects, depicts the state space, and is the action space, as presented in Table 1. As demonstrated in Figure 5, it consists of two modules: Explore and Exploit, both leveraging Chain-of-Thought (CoT) reasoning to adapt to new environments and optimize learned navigation execution. The Heuristic Process is expressed as:
Explore. The Explore module is activated when the Demand Matching module fails to identify a suitable target object within the agent’s current field of view. In this mode, the system defaults to a broad, information-seeking search of the environment. It generates exploratory actions by leveraging the VLM to interpret its visual perspective and reason about promising, unvisited areas. The objective is to efficiently survey the surroundings to uncover previously overlooked objects or navigational paths. As illustrated in Figure 2, this process begins with the system generating a concise Scene Description D from the current viewpoint image . Subsequently, it engages in explicit Reasoning R to formulate a strategy for exploration. Based on this reasoning, the system makes a Decision S, which comprises a sequence of actions designed to navigate toward unvisited regions. To prevent inefficient behavior, the module incorporates a short-term memory of prior actions and recent rotations to minimize redundant movements. Furthermore, if an obstacle impedes progress, a flag is set to true, prompting the system to dynamically replan and generate new exploratory actions, which ensures continuous progress without getting stuck in loops or depleting resources.
@@
Exploit. Once the Demand Matching module successfully identifies a target object, the system transitions from exploration to exploitation. The Exploit module shifts the agent’s strategy from a broad, open-ended search to precise, goal-directed action. This phase leverages a version of the VLM that has been fine-tuned on prior successful experiences stored in the knowledge base, optimizing it for goal achievement. This fine-tuning enhances the model’s ability to make accurate predictions and decisions based on subtle environmental cues, enabling the system to execute complex navigation tasks with high precision. Guided by a Chain-of-Thought(CoT) process, the system first generates a Scene Description D based on the current viewpoint image and the information about the target objects . It then performs focused Reasoning R to infer the most direct course of action toward the target, given D and . Finally, it produces the most optimal Decision S, a single, high-confidence action focused on progressing toward the target. By concentrating on a single, decisive action rather than a sequence of potential actions, the system improves its decisiveness and computational efficiency, ensuring rapid and effective goal attainment.
3.5. Analytic Process
The Analytic Process serves as the deliberative reasoning component of our framework, designed to reflect upon and learn from navigation failures, as illustrated in Figure 6. Through comprehensive pre-training on diverse, large-scale datasets, VLMs naturally accumulate a vast repository of world knowledge. This enables them to address complex challenges with sophisticated reasoning and insight, which is a capability that aligns perfectly with the objectives of the Analytic Process. This process moves beyond the rapid, intuitive responses of the Heuristic Process and instead relies on detailed causal analysis and deep contextual understanding to draw sound, generalizable inferences from navigational errors encountered in indoor environments.
Specifically, within the closed-loop operation of our system, any obstruction or failure encountered during the execution of the Heuristic Process triggers this reflective mechanism. When triggered, the Analytic Process utilizes the extensive world knowledge embedded within the VLMs to diagnose the situation. To do this, it ingests a snapshot of the state preceding the failure, including the target object’s location, the scene description, and the specific reasoning and decision that led to the error. The system then carefully analyzes the root cause of the event, identifies the flaw in the previous logic, and generates a corrected chain of reasoning along with a revised decision. The valuable insights gained from this reflective process, a successful pairing of a problem state with a corrected solution, are subsequently integrated into the knowledge base. This integration establishes an iterative learning loop, enabling the system to learn from its failures continuously and progressively enhance its decision-making capabilities. This virtuous cycle ultimately leads to more informed and accurate navigation strategies for future tasks and directly improves the performance of the Heuristic Process over time.